dear sir/madam,how to solve this problem?please help.

The number of positive integral solutions of XYZT=210 are


A)256


b)64


c)24


d)120


Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear h v

factorise 210 = 7*5*3*2

XYZT =210

7 can be assign to any X,Y,Z,T in = 4 ways

similerly    5 can be assign to any X,Y,Z,T in = 4 ways

              3 can be assign to any X,Y,Z,T in = 4 ways

              2 can be assign to any X,Y,Z,T in = 4 ways

 so total ways  = 4*4*4*4 =256

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

h v
13 years ago

thanks for the answer